# Simple and Provable Scaling Laws for the Test-Time Compute of Large Language Models

### v4

@@ 4.1 Results for the knockout-style algorithm
Figure 2 includes results for majority voting, a strong baseline widely adopted in practice. It is observed that, for all LLM backends (except for Llama3.1), the knockout-style algorithm consistently achieves higher accuracy when given the same number $N$ of initial candidates. Caution should be taken here: recall from Section 2.2 that the knockout-style algorithm takes $(K+1)\times N$ LLM calls for solving one problem, i.e., $5\times N$ for Llama3.1/ Qwen2.5/ Mixed and $3\times N$ for GPT-4o/ QwQ-32B, whereas majority voting only requires $N$ LLM calls. Nonetheless, the knockout-style algorithm still has advantage when this is taken into account, e.g., its accuracy at $N=8$ (resp. 4) is higher than that of majority voting at $N=64$ (resp. 16) for Mixed (resp. QwQ-32B). Moreover, based on the trends shown in Figure 2, it is most likely that for majority voting, further increasing $N$ will bring limited performance gains [3] and result in a converged accuracy lower than what can be achieved by the knockout-style algorithm.
Figure 2: Accuracy versus the number of initial candidates $N$ for the knockout-style algorithm (KO), as well as for majority voting (MV), a strong baseline widely adopted in practice.
@@
The results in Figure 2 are indeed consistent with the theorems developed in Section 2.1, which guarantee that the knockout-style algorithm can achieve an arbitrarily high success probability for any input problem satisfying Assumption 1, namely $p_{\text{gen}}>0$ and $p_{\text{comp}}>0.5$. For a problem that does not, it is still possible that the algorithm has a chance of solving it correctly (since Assumption 1 is a sufficient condition for its success and might not be necessary), but there is no formal guarantee. Consequently, for a benchmark or a distribution of input problems, denoted by $\mathcal{D}$, our algorithm is guaranteed to achieve accuracy at least ${\mathbb{P}}_{x\sim\mathcal{D}}(x\text{ satisfies the assumption})$ as its test-time compute grows. Indeed, if a benchmark contains an extremely difficult problem, e.g., “solve the P versus NP problem”, then any test-time scaling method will fail to achieve 100% accuracy on such a benchmark.
@@
Figure 3: Left: the distribution of GPQA problems, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ that are estimated with the knockout-style algorithm (Mixed). Each problem is represented by a circle if it is solved correctly at $N=64$, and by a cross otherwise. Right: accuracy versus $N$ for the knockout-style algorithm, on a filtered subset of problems satisfying $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{p}_{\text{comp}}>0.5$. The values of accuracy are calculated with new trials of the algorithm, thus statistically independent of $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$.
@@
To further bridge the empirical results with theories, let us start by estimating the parameters $p_{\text{gen}}$ and $p_{\text{comp}}$ in Assumption 1. For each problem, we define $\widehat{p}_{\text{gen}}$ as the proportion of the $N=64$ initial candidate solutions with a correct final answer, which serves as a good proxy for $p_{\text{gen}}$. To find a proxy for $p_{\text{comp}}$, we define $\widehat{p}_{\text{comp}}$ by picking all LLM calls for comparing a pair of correct and incorrect solutions throughout the knockout tournament, putting higher weights on the comparison results from later rounds of the tournament, and taking the sum of the weights of comparisons that identify the right winners [^3] The rationale for weighting the comparison results is explained as follows. In the early rounds of the knockout tournament, the comparison result between a correct-but-weak candidate and an incorrect candidate can cause a negative bias in estimating $p_{\text{comp}}$; similarly, a correct candidate might have a very high win rate against an opponent that is not only incorrect but also very weak, which can cause a positive bias in estimating $p_{\text{comp}}$. In contrast, the correct or incorrect candidates that survive the early rounds of the knockout tournament tend to be stronger ones, which make the comparison results among them (in later rounds of the tournament) more reliable and meaningful for the purpose of estimating $p_{\text{comp}}$. .
@@
Figure 7 in Appendix C.2 characterizes the distribution of GPQA and MMLU-Pro-S problems in terms of $\widehat{p}_{\text{gen}}$ (the X-axis) and $\widehat{p}_{\text{comp}}$ (the Y-axis); one such plot can also be found in Figure 3 (left). On the top half of each scatter plot are problems with $\widehat{p}_{\text{comp}}>0.5$, most of which are solved correctly by the knockout-style algorithm and represented as circles. These include some problems with small $\widehat{p}_{\text{gen}}$, for which the knockout stage successfully identifies a correct candidate even though the initial candidates are mostly incorrect. We further observe that, compared to Llama3.1 and Qwen2.5, the Mixed option achieves $\widehat{p}_{\text{gen}}>0$ and $\widehat{p}_{\text{comp}}>0.5$ for a larger proportion of problems, which explains its superior accuracy shown in Figure 2. To further consolidate this analysis, we pay special attention to the subset of problems satisfying $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{p}_{\text{comp}}>0.5$. These are approximations for the conditions stated in Assumption 1, except that those easy problems with $\widehat{p}_{\text{gen}}=1$ are excluded. We run new, independent trials of the knockout-style algorithm on this subset. Figure 3 (right) confirms that significant improvements in accuracy (e.g., 25 points for Mixed) can be achieved for such problems by scaling up $N$, which matches what our theories predict.
Intuitions: when does Assumption 1 hold?
@@
Figure 4: Accuracy versus $N$ for the knockout-style algorithm on two categories of MMLU-Pro-S.
@@ 4.2 Results for the league-style algorithm
Figure 5 shows that the accuracy of the league-style algorithm grows with $N$ for all of Llama3.1, Qwen2.5 and Mixed options on GPQA, e.g., accuracy of Mixed improves by 8 points (from $45\%$ to $53\%$) as $N$ grows to 16. The Mixed option consistently outperforms Llama3.1 and Qwen2.5, similar to the case for the knockout-style algorithm. Given the same initial candidates, the league-style algorithm achieves higher accuracy than the knockout-style algorithm does in some cases and lower in other cases, although the differences are minor in general.
@@
Let us start by finding proxies for the $p_{\text{cs}}$ and $\Delta$ parameters in Assumption 2. For the former, we define $\widehat{p}_{\text{cs}}$ as the fraction of initial candidate solutions with the correct final answer. For the latter, we define $\widehat{\Delta}\coloneqq\widehat{\mu}_{i_{1}}-\widehat{\mu}_{i_{2}}$, where $i_{1}\coloneqq\arg\max_{i\in[N]:\ y_{i}\text{ is correct}}\widehat{\mu}_{i}$ and $i_{2}\coloneqq\arg\max_{i\in[N]:\ y_{i}\text{ is incorrect}}\widehat{\mu}_{i}$ are the indices for the strongest correct candidate and for the strongest incorrect candidate, respectively. Note that by definition, for any problem with $\widehat{p}_{\text{cs}}\notin\{0,1\}$, the league-style algorithm returns a correct solution to the problem if and only if $\widehat{\Delta}>0$.
Figure 15 in Appendix C.3 characterizes the distribution of GPQA and MMLU-Pro-S problems in terms of $\widehat{p}_{\text{cs}}$ (X-axis) and $\widehat{\Delta}$ (Y-axis); the plot corresponding to GPQA and Mixed can also be found in Figure 5. It is noteworthy that there exists a non-trivial proportion of problems for which $\widehat{p}_{\text{cs}}$ is fairly small (i.e., most of the initial candidates are incorrect), yet the proposed league-style aggregation stage still manages to attain a positive $\widehat{\Delta}$ and thus identify a correct candidate for the final output. Figure 5 also confirms that significant improvements in accuracy, e.g., a 25% increase for Mixed, can be achieved on the subset of problems that approximately satisfy Assumption 2.
@@
While all previous experiments consider the round-robin version of the league-style algorithm, we also wonder if it is feasible to improve its computational efficiency by estimating the average win rate of each candidate with $M<N-1$ subsampled opponents, while maintaining its accuracy. The empirical results in Figure 5 provide a positive answer and match what our theories in Section 3 predict: (1) accuracy initially increases with $M$, which confirms the benefits of comparing each candidate with multiple opponents; (2) once $M$ exceeds a threshold (around 4 or 5) that is much smaller than $N=16$, accuracy saturates around the level achieved by the round-robin version, but at a lower computational cost.
Figure 5: Empirical results for the league-style algorithm. (1) Accuracy versus the number of initial candidates $N$ for the league-style (LG, solid lines) and knockout-style (KO, dotted lines) algorithms, given the same initial candidates. (2) The distribution of GPQA problems, characterized by $\widehat{p}_{\text{cs}}$ and $\widehat{\Delta}$ that are estimated with the Mixed option. Each problem is represented by a circle if it is solved correctly at $N=16$, and by a cross otherwise. (3) Accuracy versus $N$ on the subset of problems satisfying $10<\widehat{p}_{\text{cs}}<$ 1 and $\widehat{\Delta}>0$. The values of accuracy are calculated with new trials of the algorithm, thus statistically independent of $\widehat{p}_{\text{cs}}$ and $\widehat{\Delta}$. (4) Accuracy versus $M$, the number of subsampled opponents for each candidate, for the league-style algorithm with $N=16$.
@@ 5 Related works
There exist other test-time strategies that enjoy provable inference scaling laws in the sense of Definition 1. One example is majority voting, whose provable success requires two assumptions [3, 48]: (1) it is feasible to divide the candidate solutions into several groups and have a meaningful count for each group (which is not the case in tasks like open-ended writing, where all candidate solutions are distinct), and (2) the probability that one LLM call generates a solution belonging to the correct group is higher than that for any other group. In comparison, our proposed algorithms are free from the first restriction, and only require $p_{\text{gen}}>0$ while making additional assumptions about LLMs’ capabilities in pairwise comparisons. Another example is Best-of-N (BoN) sampling, for which deriving a provable scaling law is straightforward provided that a perfect verifier is available: if one LLM call generates a correct solution with probability $p_{\text{gen}}>0$, then the failure probability of BoN is $(1-p_{\text{gen}})^{N}$. One obvious limitation is that verifiers are unavailable or imperfect in many practical scenarios, which can hinder the performance of BoN [5, 37, 2]. We refrain from comparing our methods with BoN in our experiments, since introducing an external verifier or reward model will bring extra variability that makes it difficult to conduct a fair and meaningful empirical comparison.
Our algorithm design has drawn inspiration from various areas. For example, the essential idea of pairwise comparison has been prominent in LLM alignment [1, 30, 33] and the LLM-as-a-judge paradigm [57, 20]. Although it is possible to verify, score or refine a solution by itself [14, 25, 6, 11], it is often much easier (for LLMs or human) to detect the errors or hallucinations in an incorrect solution when it is placed right next to a correct one, or evaluate the quality of a solution by comparing it to another one. The knockout and league tournaments have also been investigated in prior LLM research [17, 23, 56, 19, 13], albeit with purposes or implementations that are different from ours. Given this context, we remark that the main novelty and contributions in this work are perhaps less about the proposed two-stage algorithms themselves, but rather more about developing rigorous and theoretical understanding of their underlying assumptions and efficacy (via clearly identifying sufficient conditions for boosting their success probability up to $100\%$ and formally deriving quantitative bounds for their computational and sample complexities), and demonstrating their promising empirical performance through extensive experiments.

### v5

@@ 4.1 Results for the knockout-style algorithm
Figure 2 includes results for majority voting, a strong baseline widely adopted in practice. It is observed that, for all LLM backends (except for Llama3.1), the knockout-style algorithm consistently achieves higher accuracy when given the same number $N$ of initial candidates. Caution should be taken here: recall from Section 2.2 that the knockout-style algorithm takes $(K+1)\times N$ LLM calls for solving one problem, i.e., $5\times N$ for Llama3.1/ Qwen2.5/ Mixed and $3\times N$ for GPT-4o/ QwQ-32B, whereas majority voting only requires $N$ LLM calls. Nonetheless, the knockout-style algorithm still has advantage when this is taken into account, e.g., its accuracy at $N=8$ (resp. 4) is higher than that of majority voting at $N=64$ (resp. 16) for Mixed (resp. QwQ-32B). Moreover, based on the trends shown in Figure 2, it is most likely that for majority voting, further increasing $N$ will bring limited performance gains chen2024llm and result in a converged accuracy lower than what can be achieved by the knockout-style algorithm.
Refer to captionRefer to caption
Figure 2: Accuracy versus the number of initial candidates NN for the knockout-style algorithm (KO), as well as for majority voting (MV), a strong baseline widely adopted in practice.
@@
The results in Figure 2 are indeed consistent with the theorems developed in Section 2.1, which guarantee that the knockout-style algorithm can achieve an arbitrarily high success probability for any input problem satisfying Assumption 2.1, namely $p_{\text{gen}}>0$ and $p_{\text{comp}}>0.5$. For a problem that does not, it is still possible that the algorithm has a chance of solving it correctly (since Assumption 2.1 is a sufficient condition for its success and might not be necessary), but there is no formal guarantee. Consequently, for a benchmark or a distribution of input problems, denoted by $\mathcal{D}$, our algorithm is guaranteed to achieve accuracy at least ${\mathbb{P}}_{x\sim\mathcal{D}}(x\text{ satisfies the assumption})$ as its test-time compute grows. Indeed, if a benchmark contains an extremely difficult problem, e.g., “solve the P versus NP problem”, then any test-time scaling method will fail to achieve 100% accuracy on such a benchmark.
@@
Figure 3: Left: the distribution of GPQA problems, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ that are estimated with the knockout-style algorithm (Mixed). Each problem is represented by a circle if it is solved correctly at $N=64$, and by a cross otherwise. Right: accuracy versus NN for the knockout-style algorithm and majority voting (both with Mixed), on a filtered subset of problems satisfying $0<\widehat{p}_{\text{gen}}<1$ and $\widehat{p}_{\text{comp}}>\tau$, where $\tau\in\{0.5,0.6,0.7\}$. The values of accuracy are calculated with new trials of the algorithm, thus statistically independent of $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$.
@@
To further bridge the empirical results with theories, let us start by estimating the parameters $p_{\text{gen}}$ and $p_{\text{comp}}$ in Assumption 2.1. For each problem, we define $\widehat{p}_{\text{gen}}$ as the proportion of the $N=64$ initial candidate solutions with a correct final answer, which serves as a good proxy for $p_{\text{gen}}$. To find a proxy for $p_{\text{comp}}$, we define $\widehat{p}_{\text{comp}}$ by picking all LLM calls for comparing a pair of correct and incorrect solutions throughout the knockout tournament, putting higher weights on the comparison results from later rounds of the tournament, and taking the sum of the weights of comparisons that identify the right winners [^4] The rationale for weighting the comparison results is explained as follows. In the early rounds of the knockout tournament, the comparison result between a correct-but-weak candidate and an incorrect candidate can cause a negative bias in estimating $p_{\text{comp}}$; similarly, a correct candidate might have a very high win rate against an opponent that is not only incorrect but also very weak, which can cause a positive bias in estimating $p_{\text{comp}}$. In contrast, the correct or incorrect candidates that survive the early rounds of the knockout tournament tend to be stronger ones, which make the comparison results among them (in later rounds of the tournament) more reliable and meaningful for the purpose of estimating $p_{\text{comp}}$. .
@@
Figure 7 in Appendix D.2 characterizes the distribution of GPQA and MMLU-Pro-S problems in terms of $\widehat{p}_{\text{gen}}$ (the X-axis) and $\widehat{p}_{\text{comp}}$ (the Y-axis); one such plot can also be found in Figure 3 (left). On the top half of each scatter plot are problems with $\widehat{p}_{\text{comp}}>0.5$, most of which are solved correctly by the knockout-style algorithm and represented as circles. These include some problems with small $\widehat{p}_{\text{gen}}$, for which the knockout stage successfully identifies a correct candidate even though the initial candidates are mostly incorrect. We further observe from Figure 7 that, compared to Llama3.1 and Qwen2.5, the Mixed option achieves $\widehat{p}_{\text{gen}}>0$ and $\widehat{p}_{\text{comp}}>0.5$ for a larger proportion of problems, which explains its superior accuracy shown in Figure 2. To further consolidate this analysis, we pay special attention to the subset of problems satisfying $0<\widehat{p}_{\text{gen}}<1$ and $\widehat{p}_{\text{comp}}>0.5$. These are approximations for the conditions stated in Assumption 2.1, except that those easy problems with $\widehat{p}_{\text{gen}}=1$ are excluded. We run new, independent trials of the knockout-style algorithm (Mixed) on this subset. Figure 3 (right) confirms that significant improvements in accuracy (from $55\%$ to $80\%$) can be achieved by scaling up $N$, which matches what our theories predict. Unsurprisingly, the scaling curve still plateaus eventually (since $\widehat{p}_{\text{comp}}>0.5$ is merely a proxy for $p_{\text{comp}}>0.5$), and tightening the filtering condition (e.g., $\widehat{p}_{\text{comp}}>0.6$ or $0.7$) will bring it closer to $100\%$ accuracy.
Intuitions: when does Assumption 2.1 hold?
@@
Figure 4: Accuracy versus NN for the knockout-style algorithm on two categories of MMLU-Pro-S.
@@ 4.2 Results for the league-style algorithm
Figure 5 (a) shows that the accuracy of the league-style algorithm grows with $N$ for all of Llama3.1, Qwen2.5 and Mixed options on GPQA, e.g., accuracy of Mixed improves by 8 points (from $45\%$ to $53\%$) as $N$ grows to 16. The Mixed option consistently outperforms Llama3.1 and Qwen2.5, similar to the case for the knockout-style algorithm. Given the same initial candidates, the league-style algorithm achieves higher accuracy than the knockout-style algorithm does in some cases and lower in other cases, although the differences are minor in general.
@@
Let us start by finding proxies for the $p_{\text{cs}}$ and $\Delta$ parameters in Assumption 3.1. For the former, we define $\widehat{p}_{\text{cs}}$ as the fraction of initial candidate solutions with the correct final answer. For the latter, we define $\widehat{\Delta}\coloneqq\widehat{\mu}_{i_{1}}-\widehat{\mu}_{i_{2}}$, where $i_{1}\coloneqq\arg\max_{i\in[N]:\ y_{i}\text{ is correct}}\widehat{\mu}_{i}$ and $i_{2}\coloneqq\arg\max_{i\in[N]:\ y_{i}\text{ is incorrect}}\widehat{\mu}_{i}$ are the indices for the strongest correct candidate and for the strongest incorrect candidate, respectively. Note that by definition, for any problem with $\widehat{p}_{\text{cs}}\notin\{0,1\}$, the league-style algorithm returns a correct solution to the problem if and only if $\widehat{\Delta}>0$.
Figure 15 in Appendix D.3 characterizes the distribution of GPQA and MMLU-Pro-S problems in terms of $\widehat{p}_{\text{cs}}$ (X-axis) and $\widehat{\Delta}$ (Y-axis); the plot corresponding to GPQA and Mixed can also be found in Figure 5 (b). It is noteworthy that there exists a non-trivial proportion of problems for which $\widehat{p}_{\text{cs}}$ is fairly small (i.e., most of the initial candidates are incorrect), yet the proposed league-style aggregation stage still manages to attain a positive $\widehat{\Delta}$ and thus identify a correct candidate for the final output. On the other hand, for problems with $\widehat{\Delta}<0$ (which indicates violations of Assumption 3.1), there is no success guarantee for the algorithm. Figure 5 (c) further confirms that significant improvements in accuracy, e.g., a 25% increase for Mixed, can be achieved on the subset of problems that approximately satisfy Assumption 3.1.
@@
While all previous experiments consider the round-robin version of the league-style algorithm, we also wonder if it is feasible to improve its computational efficiency by estimating the average win rate of each candidate with M<N1M<N-1 subsampled opponents, while maintaining its accuracy. The empirical results in Figure 5 (d) provide a positive answer and match what our theories in Section 3 predict: (1) accuracy initially increases with $M$, which confirms the benefits of comparing each candidate with multiple opponents; (2) once $M$ exceeds a threshold (around 4 or 5) that is much smaller than $N=16$, accuracy saturates around the level achieved by the round-robin version, but at a lower computational cost.
Figure 5: Empirical results for the league-style algorithm. (a) Accuracy versus the number of initial candidates NN for the league-style (LG, solid lines) and knockout-style (KO, dotted lines) algorithms, given the same initial candidates. (b) The distribution of GPQA problems, characterized by $\widehat{p}_{\text{cs}}$ and $\widehat{\Delta}$ that are estimated with the Mixed option. Each problem is represented by a circle if it is solved correctly at $N=16$, and by a cross otherwise. (c) Accuracy versus NN on the subset of problems satisfying $0<\widehat{p}_{\text{cs}}<1$ and $\widehat{\Delta}>0$. The values of accuracy are calculated with new trials of the algorithm, thus statistically independent of $\widehat{p}_{\text{cs}}$ and $\widehat{\Delta}$. (d) Accuracy versus $M$, the number of subsampled opponents for each candidate, for the league-style algorithm with $N=16$.
@@ 5 Related works
There exist other test-time strategies that enjoy provable inference scaling laws in the sense of Definition 1.1. One example is majority voting, whose provable success requires two assumptions chen2024llm ; wu2024empiricalanalysiscomputeoptimalinference : (1) it is feasible to divide the candidate solutions into several groups and have a meaningful count for each group (which is not the case in tasks like open-ended writing, where all candidate solutions are distinct), and (2) the probability that one LLM call generates a solution belonging to the correct group is higher than that for any other group. In comparison, our proposed algorithms are free from the first restriction, and only require $p_{\text{gen}}>0$ while making additional assumptions about LLMs’ capabilities in pairwise comparisons. Another example is best-of-N (BoN) sampling, for which deriving a provable scaling law is straightforward provided that a perfect verifier is available: if one LLM call generates a correct solution with probability $p_{\text{gen}}>0$, then the failure probability of BoN is $(1-p_{\text{gen}})^{N}$. One obvious limitation is that verifiers are unavailable or imperfect in many practical scenarios, which can hinder the performance of BoN cobbe2021trainingverifierssolvemath ; stroebl2024inferencescalingflawslimits ; brown2024largelanguagemonkeysscaling . We refrain from comparing our methods with BoN in our experiments, since introducing an external verifier or reward model will bring extra variability that makes it difficult to conduct a fair and meaningful empirical comparison.
Our algorithm design has drawn inspiration from various areas. For example, the essential idea of pairwise comparison has been prominent in LLM alignment Bradley1952RankAO ; Ouyang2022TrainingLM ; rafailov2023direct and the LLM-as-a-judge paradigm zheng2023judging ; li2024generationjudgmentopportunitieschallenges . Although it is possible to verify, score or refine a solution by itself kadavath2022languagemodelsmostlyknow ; madaan2023selfrefine ; davis2024networksnetworkscomplexityclass ; huang2024large , it is often much easier (for LLMs or human) to detect the errors or hallucinations in an incorrect solution when it is placed right next to a correct one, or evaluate the quality of a solution by comparing it to another one. The knockout and league tournaments have also been investigated in prior LLM research kawabata2024rationaleawareanswerverificationpairwise ; liu2024statistical ; zhao2023slichfsequencelikelihoodcalibration ; lee2024aligning ; jiang2023llmblender , albeit with purposes or implementations that are different from ours. Given this context, we remark that the main novelty and contributions in this work are perhaps less about the proposed two-stage algorithms themselves, but rather more about developing rigorous and theoretical understanding of their underlying assumptions and efficacy (via clearly identifying sufficient conditions for boosting their success probability up to $100\%$ and formally deriving quantitative bounds for their computational and sample complexities), and demonstrating their promising empirical performance through extensive experiments.